Write the first four terms of the sequence. Simplify your answer

a n = 3n/3n+3

a1=

a2=

a3=

a4=
Answer #1

an = 3n/3n+3

a1 = 3*(1)/3*(1)+3 = 4

a2 =(3*(2))/(3*(2))+3 = 4

a3 = (3*(3))/(3*(3))+3 = 4

a4 = (3*(4))/(3*(4))+3 = 4